# Optimizing Attention with Mirror Descent: Generalized Max-Margin Token Selection

### v3

@@ 3.1 Optimizing Attention with Fixed Head $v$
$\begin{cases}\left[W(k+1)\right]_{ij}\leftarrow\left|[W(k)]^{+}_{ij}\right|^{\frac{1}{p-1}}\cdot\textnormal{sign}\left([W(k)]^{+}_{ij}\right),\vspace{.3cm}\\ [W(k)]^{+}_{ij}:=|[W(k)]_{ij}|^{p-1}\textnormal{sign}([W(k)]_{ij})-\eta[\nabla\mathcal{L}(W(k))]_{ij}.\end{cases}$ ($\ell_{p}$-AttGD)
@@
$\|W(k+1)\|_{p,p}-\|W(k)\|_{p,p}\leq\mathcal{O}\big{(}\|W(k)\|_{p,p}^{2-p}\big{)},\quad\textnormal{for all}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ k=0,1,2,\dots\leavevmode\nobreak\ .$
@@
[…]
@@
[…]
@@
$W^{(R)}:=\operatorname*{arg\,min}_{W\in\mathbb{R}^{d\times d}}\leavevmode\nobreak\ \mathcal{L}(W)\quad\textnormal{subj. to}\quad\|W\|_{p,p}\leq R.$ ($\ell_{p}$-AttRP)
@@
$S_{p,\mu}(W)$ :=$:=\left\{W^{\prime}\in\mathbb{R}^{d\times d}\mid D_{\psi}\left(\frac{W}{\|W\|_{p,p}},\frac{W^{\prime}}{\|W^{\prime}\|_{p,p}}\right)\leq\mu\right\},$ (4a)
@@
These theorems show that as the parameters grow large enough and approach a locally optimal direction, they will keep moving toward that direction.
@@
$D_{\psi}\left(\frac{W^{\alpha}_{\mathrm{mm}}}{\|W^{\alpha}_{\mathrm{mm}}\|_{p,p}},\frac{W(k)}{\|W(k)\|_{p,p}}\right)=\mathcal{O}\left(\begin{cases}\frac{\log\log k}{\log k}&\textnormal{if}\leavevmode\nobreak\ \leavevmode\nobreak\ p>2,\\ \frac{(\log\log k)^{2}}{\log k}&\textnormal{if}\leavevmode\nobreak\ \leavevmode\nobreak\ p=2,\\ \frac{1}{(\log k)^{p-1}}&\textnormal{otherwise}.\end{cases}\right).$ (5)
Note that, in the left-hand side of (5), there is a dependence on $p$ in the Bregman divergence $D_{\psi}$ itself as well. Despite optimizing a highly nonlinear, nonconvex softmax function, we achieve a convergence rate similar to GD in linear binary classification (Ji and Telgarsky, 2018, Theorem 1.1) (up to a $\log\log k$ factor).
@@
Since we aim to show the parameter converges in direction to the cone center, $W_{\mathrm{mm}}^{\alpha}$, we need conditions ensuring the parameters remain in the cone. We formalize this in Lemma 28 and prove that for any $\mu>0$ and locally optimal tokens $(\alpha_{i})_{i=1}^{n}$ (Definition 3), there exist constants $R,\mu^{\prime}>0$ depending on the dataset and $\mu$, such that if $W(0)\in C_{p,\mu^{\prime},R}(W^{\alpha}_{\mathrm{mm}})$, then $W(k)\in C_{p,\mu,R}(W^{\alpha}_{\mathrm{mm}})$ for all $k$, meaning the iterates remain within a larger cone; see Figure 2.
@@
$|W(k+1)\|_{p,p}^{p-1}$ $\geq\|W(k)\|_{p,p}^{p-1}$
$+\frac{\eta}{\|W(k)\|_{p,p}}\langle-\nabla\mathcal{L}(W(k)),W(k)\rangle.$
@@ 3.2 Training Dynamics of Mirror Descent for Joint Optimization of $W$ and $v$
$\begin{cases}\left[W(k+1)\right]_{ij}\leftarrow\left|[W(k)]^{+}_{ij}\right|^{\frac{1}{p-1}}\cdot\textnormal{sign}\left([W(k)]^{+}_{ij}\right),\vspace{.3cm}\\ [W(k)]^{+}_{ij}:=|[W(k)]_{ij}|^{p-1}\textnormal{sign}([W(k)]_{ij})-\eta_{W}[\nabla_{W}\mathcal{L}(W(k),v(k))]_{ij},\vspace{.3cm}\\ [v(k+1)]_{i}\leftarrow\left|[v(k)]_{i}^{+}\right|^{\frac{1}{p-1}}\cdot\textnormal{sign}([v(k)]_{i}^{+}),\vspace{.3cm}\\ [v(k)]_{i}^{+}:=|[v(k)]_{i}|^{p-1}\textnormal{sign}([v(k)]_{i})-\eta_{v}[\nabla_{v}\mathcal{L}(W(k),v(k))]_{i}.\end{cases}$ ($\ell_{p}$-JointGD)
@@
$\begin{split}v_{\mathrm{mm}}&:=\operatorname*{arg\,min}_{v\in\mathbb{R}^{d}}\|v\|_{p}\\ &\textnormal{subj.\leavevmode\nobreak\ to}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ y_{i}X_{i\alpha_{i}}^{\top}v\geq 1,\leavevmode\nobreak\ \leavevmode\nobreak\ \textnormal{for all }\leavevmode\nobreak\ i\in[n].\end{split}$ ($\ell_{p}$-SVM)
@@
Figure 4: Average directional Bregman divergence between the (a) $\ell_{1.75}$, (b) $\ell_{2}$, and (c) $\ell_{3}$ optimization paths and the ($\ell_{p}$-AttSVM) solutions for $p=1.75,2$, and $3$ at each training iteration from 100 trials. The shaded area represents the standard deviation of the directional Bregman divergence.
@@
4 Experimental Results
4.1 Synthetic Data Experiments
4.1.1 $\ell_{p}$-AttGD Experiment
Figure 4 shows the directional Bregman divergence between the ($\ell_{p}$-AttSVM) solution and the $\ell_{q}$ optimization path for each pair $p,q\in\{1.75,2,3\}$. In Figure 4(a), the divergence converges to 00 only for the ($\ell_{p}$-AttSVM) ($p=1.75$) solution, indicating that the $\ell_{1.75}$ path does not converge to the $p=2$ or $3$ solutions. The shrinking standard deviation shows consistent behavior. Similarly, Figures 4(b) and 4(c) show the divergence converging to 00 for the corresponding ($\ell_{p}$-AttSVM) solution, demonstrating that the $\ell_{p}$ optimization path converges to the ($\ell_{p}$-AttSVM) solution, with the direction of convergence changing with $p$. In addition to the directional Bregman divergence, we can also observe the convergence in direction for one trial (initialized at zero) directly by plotting how two of the entries of $W$ change during training simultaneously and plotting it on a Cartesian graph, then showing that the path it follows converges to the direction of the ($\ell_{p}$-AttSVM) solution. As we can see in Figure 5, each of the $\ell_{p}$ optimization paths follows the direction of the corresponding ($\ell_{p}$-AttSVM) solution.
@@ 4.1.2 $\ell_{p}$-JointGD Experiments
We use the data from Example  2 in Appendix  G.3 to train a model using  $\ell_{p}$-JointGD for $p=1.75,2,$ and $3$. The comparison between the iterates and the SVM solutions in Figure 6 shows that the iterates of $W$ and $v$ converge to the  $\ell_{p}$-AttSVM and  $\ell_{p}$-SVM directions, respectively, for each of $p=1.75,2,$ and $3$. These convergence are similar to Theorem 31, as in both this experiment and that theorem, we get that the iterates converge to the SVM problem solutions. In addition to these iterates, we record the evolution of the average softmax probability of the optimal token, along with the average logistic probability of the model, which we define to be $1/n\sum_{i=1}^{n}1/(1+\exp(-\gamma_{i}a_{i}))$.
@@ 4.2 Real Data Experiments
This section presents evidence of improved generalization, token selection, and explainability from training an attention network with MD instead of other algorithms such as GD and Adam (Kingma and Ba, 2014). Specifically, we compare the generalization, token selection, and model weight distribution of the MD and GD algorithms in training a transformer classification model on the Stanford Large Movie Review Dataset; and the generalization and model weight distribution of the MD and Adam in training a vision transformer (ViT) model (Dosovitskiy et al., 2020) on CIFAR-10.
@@ 4.2.1 Comparison with Gradient Descent
[…]
Figure 8: The attention map generated by the resulting models that were trained using $\ell_{1.1}$ and GD for five sample sentences. For three out of five of the sample sentences, the model trained using $\ell_{1.1}$ selects the optimal token better than the model trained using GD.
To investigate this further, we look at how the attention layers of the model select the tokens from simple reviews that GPT-4o generated and investigate how much the attention layer focuses on a particular token that truly determines whether the whole review was a positive one or a negative one. We chose these pivotal tokens using GPT-4o as well. We do this procedure to the model trained using $\ell_{1.1}$ and the GD and tabulate the full results in Appendix G (we provide five of the results in Figure 8). We can see that the $\ell_{1.1}$ also outperforms the GD in token selection.
@@
Finally, we collect the training weights from the resulting models trained by $\ell_{1.1}$ and the GD and plot a histogram of their absolute values in Figure 9. Specifically, we take the histogram of the components of the key, query, and value matrices. The figures show that the resulting model that was trained using $\ell_{1.1}$ is sparser than the one trained using GD, which could hint at a potential explanation as to why $\ell_{1.1}$ can outperform the standard GD algorithm when it is used to train attention-based models.
@@ 4.2.2 Comparison with Adam
Figure 10: The test accuracy of training a ViT network using $\ell_{1.1}$-MD and Adam.
We compare the $\ell_{1.1}$ algorithm with the Adam algorithm in training a ViT model on the CIFAR-10 dataset. We specify the ViT model architecture in Appendix G.5. The resulting test accuracies for the two algorithms are reported in Figure 10. As we can see, the two algorithms achieve a similar level of test accuracy at the end, which means that our algorithm can reach the SOTA performance.
@@
We also plot the weights’ absolute value of the resulting two models that were trained by MD and Adam in Figure 11, specifically the weights from the attention layer, which consists of the key, query, and value matrices. As in Figure 9, $\ell_{1.1}$-MD creates sparser parameter set, this time compared to the one generated by Adam, which shows a potential explanation of $\ell_{1.1}$-MD’s performance.
5 Conclusion

### v5

@@ 3.1 Optimizing Attention with Fixed Head $v$
$\begin{cases}\left[W(k+1)\right]_{ij}\leftarrow\left|[W(k)]^{+}_{ij}\right|^{\frac{1}{p-1}}\cdot\textnormal{sign}\left([W(k)]^{+}_{ij}\right),\vskip 8.5359pt\\ [W(k)]^{+}_{ij}:=|[W(k)]_{ij}|^{p-1}\textnormal{sign}([W(k)]_{ij})-\eta[\nabla\mathcal{L}(W(k))]_{ij}.\end{cases}$ ($\ell_{p}$-AttGD)
@@
$\|W(k+1)\|_{p,p}-\|W(k)\|_{p,p}\leq\mathcal{O}\big(\|W(k)\|_{p,p}^{2-p}\big),\quad\textnormal{for all}\penalty 10000\ \penalty 10000\ \penalty 10000\ k=0,1,2,\dots\penalty 10000\ .$
@@
$\textnormal{subj.\penalty 10000\ to}\penalty 10000\ \penalty 10000\ (X_{i\alpha_{i}}-X_{it})^{\top}Wz_{i}\geq 1,\penalty 10000\ \penalty 10000\ \textnormal{for all}\penalty 10000\ \penalty 10000\ t\in[T]-\{\alpha_{i}\},\penalty 10000\ i\in[n].$
@@
Consider the matrices $X_{1}=[5,\penalty 10000\ 0;\penalty 10000\ 0,\penalty 10000\ 1]$ and $X_{2}=[-5,\penalty 10000\ 0;\penalty 10000\ 0,\penalty 10000\ -1]$ with $y_{1}=-y_{2}=1$. Let $X_{i1}$ be the optimal token and $X_{it}$ be the others. Solving Problem ($\ell_{p}$-AttSVM) with $p=3$ and setting $z_{i}=X_{i1}$, we obtain the solution $W_{\mathrm{mm}}:=W^{\alpha}_{\mathrm{mm}}=[0.03846,\penalty 10000\ 0;\penalty 10000\ -0.00769,\penalty 10000\ 0]$. Figure 1 illustrates how the optimal tokens $X_{11}$ and $X_{21}$ are separated by the dashed decision boundaries. These boundaries are orthogonal to the vectors $W_{\mathrm{mm}}z_{i}$ and indicate the hyperplanes that separate the sequences based on the optimal token in each case.
@@
$W^{(R)}:=\operatorname*{arg\,min}_{W\in\mathbb{R}^{d\times d}}\penalty 10000\ \mathcal{L}(W)\quad\textnormal{subj. to}\quad\|W\|_{p,p}\leq R.$ ($\ell_{p}$-AttRP)
@@
$S_{p,\mu}(W)$ $:=\left\{W^{\prime}\in\mathbb{R}^{d\times d}\mid D_{\psi}\left(\frac{W}{\|W\|_{p,p}},\frac{W^{\prime}}{\|W^{\prime}\|_{p,p}}\right)\leq\mu\right\},$ (4a)
@@
These theorems show that as the parameters grow large enough and approach a locally optimal direction, they will keep moving toward that direction. While our results build on the max-margin token selection framework of Tarzanagh et al. (2023), the extension to $\ell_{p}$-mirror descent is nontrivial: we establish directional convergence under non-quadratic potentials where standard Euclidean arguments fail, and provide the first explicit finite-time poly-logarithmic convergence rates for attention optimization in Theorem 11.
@@
$D_{\psi}\left(\frac{W^{\alpha}_{\mathrm{mm}}}{\|W^{\alpha}_{\mathrm{mm}}\|_{p,p}},\frac{W(k)}{\|W(k)\|_{p,p}}\right)=\mathcal{O}\left(\begin{cases}\frac{\log\log k}{\log k}&\textnormal{if}\penalty 10000\ \penalty 10000\ p>2,\\ \frac{(\log\log k)^{2}}{\log k}&\textnormal{if}\penalty 10000\ \penalty 10000\ p=2,\\ \frac{1}{(\log k)^{p-1}}&\textnormal{otherwise}.\end{cases}\right).$ (5)
Establishing these rates required developing new analytical tools beyond the $\ell_{2}$ blueprint: we prove $\ell_{p,p}$-norm growth bounds under non-quadratic potentials (Lemma 4), introduce double-cone stability arguments to handle joint norm and directional constraints (Lemma 28), and work throughout with Bregman divergences rather than Euclidean geometry.
@@
Note that in the left-hand side of (5), there is a dependence on $p$ in the Bregman divergence $D_{\psi}$ itself as well. Despite optimizing a highly nonlinear, nonconvex softmax function, we achieve a convergence rate similar to GD in linear binary classification (Ji and Telgarsky, 2018, Theorem 1.1) (up to a $\log\log k$ factor).
Since we aim to show the parameter converges in direction to the cone center $W_{\mathrm{mm}}^{\alpha}$, we need conditions ensuring the parameters remain in the cone. Unlike $\ell_{2}$ gradient descent where cone invariance follows from Pythagorean identities, the $\ell_{p}$ setting requires a novel double-cone construction: we must ensure iterates simultaneously remain within the directional cone $S_{p,\mu_{0}}(W_{\mathrm{mm}}^{\alpha})$ and satisfy the norm constraint $\|W\|_{p,p}\geq R_{\pi}$—a strictly joint requirement not present in prior mirror descent analyses (Sun et al., 2022). We formalize this in Lemma 28 and prove that for any $\mu>0$ and locally optimal tokens $(\alpha_{i})_{i=1}^{n}$ (Definition 3), there exist constants $R,\mu^{\prime}>0$ depending on the dataset and $\mu$ such that if $W(0)\in C_{p,\mu^{\prime},R}(W^{\alpha}_{\mathrm{mm}})$, then $W(k)\in C_{p,\mu,R}(W^{\alpha}_{\mathrm{mm}})$ for all $k$, meaning the iterates remain within a larger cone; see Figure 2.
@@
$|W(k+1)\|_{p,p}^{p-1}$ $\geq\|W(k)\|_{p,p}^{p-1}+\frac{\eta}{\|W(k)\|_{p,p}}\langle-\nabla\mathcal{L}(W(k)),W(k)\rangle.$
@@
$\langle\nabla\mathcal{L}(W),V\rangle=\Omega\left(\textnormal{exp}\left(-\frac{\|W\|_{p,p}}{\|W^{\alpha}_{\mathrm{mm}}\|_{p,p}}\left(1+\frac{1}{2}\delta\right)\right)\right)>0.$
@@ 3.2 Training Dynamics of Mirror Descent for Joint Optimization of $W$ and $v$
$\hskip-20.0pt\begin{cases}\left[W(k+1)\right]_{ij}\leftarrow\left|[W(k)]^{+}_{ij}\right|^{\frac{1}{p-1}}\cdot\textnormal{sign}\left([W(k)]^{+}_{ij}\right),\vskip 8.5359pt\\ [W(k)]^{+}_{ij}:=|[W(k)]_{ij}|^{p-1}\textnormal{sign}([W(k)]_{ij})-\eta_{W}[\nabla_{W}\mathcal{L}(W(k),v(k))]_{ij},\vskip 8.5359pt\\ [v(k+1)]_{i}\leftarrow\left|[v(k)]_{i}^{+}\right|^{\frac{1}{p-1}}\cdot\textnormal{sign}([v(k)]_{i}^{+}),\vskip 8.5359pt\\ [v(k)]_{i}^{+}:=|[v(k)]_{i}|^{p-1}\textnormal{sign}([v(k)]_{i})-\eta_{v}[\nabla_{v}\mathcal{L}(W(k),v(k))]_{i}.\end{cases}$ ($\ell_{p}$-JointGD)
@@
$\begin{split}v_{\mathrm{mm}}&:=\operatorname*{arg\,min}_{v\in\mathbb{R}^{d}}\|v\|_{p}\\ &\textnormal{subj.\penalty 10000\ to}\penalty 10000\ \penalty 10000\ \penalty 10000\ y_{i}X_{i\alpha_{i}}^{\top}v\geq 1,\penalty 10000\ \penalty 10000\ \textnormal{for all }\penalty 10000\ i\in[n].\end{split}$ ($\ell_{p}$-SVM)
@@
4 Experimental Results
4.1 Synthetic Data Experiments
We describe the setup of the experiments for $\ell_{p}$-AttGD and $\ell_{p}$-JointGD and their results.
4.1.1 $\ell_{p}$-AttGD Experiment
@@
Figure 4: Average directional Bregman divergence between the (a) $\ell_{1.75}$, (b) $\ell_{2}$, and (c) $\ell_{3}$ optimization paths and the ($\ell_{p}$-AttSVM) solutions for $p=1.75,2$, and 33 at each training iteration from 100 trials. The shaded area represents the standard deviation of the directional Bregman divergence.
@@
Figure 4 shows the directional Bregman divergence between the ($\ell_{p}$-AttSVM) solution and the $\ell_{q}$ optimization path for each pair $p,q\in\{1.75,2,3\}$. In Figure 4(a), the divergence converges to 0 only for the ($\ell_{p}$-AttSVM) ($p=1.75$) solution, indicating that the $\ell_{1.75}$ path does not converge to the $p=2$ or $3$ solutions. The shrinking standard deviation shows consistent behavior. Similarly, Figures 4(b) and 4(c) show the divergence converging to 0 for the corresponding ($\ell_{p}$-AttSVM) solution, demonstrating that the $\ell_{p}$ optimization path converges to the ($\ell_{p}$-AttSVM) solution, with the direction of convergence changing with $p$. In addition to the directional Bregman divergence, we can also observe the convergence in direction for one trial (initialized at zero) directly by plotting how two of the entries of $W$ change during training simultaneously and plotting it on a Cartesian graph, then showing that the path it follows converges to the direction of the ($\ell_{p}$-AttSVM) solution. As we can see in Figure 5, each of the $\ell_{p}$ optimization paths follows the direction of the corresponding ($\ell_{p}$-AttSVM) solution.
@@ 4.1.2 $\ell_{p}$-JointGD Experiments
We use the data from Example 2 in Appendix G.3 to train a model using $\ell_{p}$-JointGD for $p=1.75,2,$ and $3$. The comparison between the iterates and the SVM solutions in Figure 6 shows that the iterates of $W$ and $v$ converge to the $\ell_{p}$-AttSVM and $\ell_{p}$-SVM directions, respectively, for each of $p=1.75,2,$ and $3$. These convergence are similar to Theorem 31, as in both this experiment and that theorem, we get that the iterates converge to the SVM problem solutions. In addition to these iterates, we record the evolution of the average softmax probability of the optimal token, along with the average logistic probability of the model, which we define to be $1/n\sum_{i=1}^{n}1/(1+\exp(-\gamma_{i}a_{i}))$.
@@ 4.2 Real Data Experiments
This section presents evidence of improved generalization, token selection, and explainability from training an attention network with MD instead of other algorithms such as GD and Adam (Kingma and Ba, 2014). Specifically, we compare the generalization, token selection, and model weight distribution of the MD and GD algorithms in training a transformer classification model on the Stanford Large Movie Review Dataset; and the generalization and model weight distribution of the MD and Adam in training a vision transformer (ViT) model (Dosovitskiy et al., 2021a) on CIFAR-10.
@@ 4.2.1 Comparison with Gradient Descent
We trained a transformer classification model on the Stanford Large Movie Review Dataset (Maas et al., 2011) using MD with $\ell_{1.1}$, $\ell_{2}$, and $\ell_{3}$ potentials. The models are similar to the encoder module in Vaswani et al. (2017), with the last layer being a linear classification layer on the feature representation of the first $\texttt{CLS}]$ token. We put the details of the classification model in Appendix G.4. Table 1 summarizes the resulting test accuracy of several variants of that model when trained with the three algorithms, which shows that the $\ell_{1.1}$ potential MD outperforms the other MD algorithms, including the one with the $\ell_{2}$ potential, which is equivalent to the GD.
Figure 8: The attention map generated by the resulting models that were trained using $\ell_{1.1}$MD and GD for five sample sentences. For three out of five of the sample sentences, the model trained using $\ell_{1.1}$MD selects the optimal token better than the model trained using GD.
To investigate this further, we look at how the attention layers of the model select the tokens from simple reviews that GPT-4o generated and investigate how much the attention layer focuses on a particular token that truly determines whether the whole review was a positive one or a negative one. We chose these pivotal tokens using GPT-4o as well. We do this procedure to the model trained using $\ell_{1.1}$MD and the GD and tabulate the full results in Appendix G (we provide five of the results in Figure 8). We can see that the $\ell_{1.1}$MD also outperforms the GD in token selection.
@@
Finally, we collect the training weights from the resulting models trained by $\ell_{1.1}$MD and the GD and plot a histogram of their absolute values in Figure 9. Specifically, we take the histogram of the components of the key, query, and value matrices. The figures show that the resulting model that was trained using $\ell_{1.1}$MD is sparser than the one trained using GD, which could hint at a potential explanation as to why $\ell_{1.1}$MD can outperform the standard GD algorithm when it is used to train attention-based models.
@@ 4.2.2 Comparison with Adam
We compare the $\ell_{1.1},\ell_{1.75},\ell_{2},\ell_{3}$MD, and Adam algorithms in training a ViT model on the CIFAR-10 dataset. We specify the ViT model architecture in Appendix G.5.
The resulting test accuracies for all algorithms are reported in Figure LABEL:experiment-adam. Notably, all five optimization methods—spanning the $\ell_{p}$-MD family with $p\in\{1.1,1.75,2,3\}$ and Adam—converge to approximately $82\%$ test accuracy on CIFAR-10, demonstrating that our $\ell_{p}$-MD framework achieves state-of-the-art performance competitive with widely-used adaptive optimizers. The consistent performance across different values of $p$ confirms that the beneficial sparsity and interpretability properties induced by the $\ell_{p}$ geometry (particularly for $p$ close to 1) come at no cost to generalization capability.
Figure 10: Test accuracy comparison across different optimizers on CIFAR-10 and -100.
Furthermore, we examine the learned weight distributions by plotting the absolute values of attention layer parameters (key, query, and value matrices) from models trained with different optimizers in Figure 11. Consistent with our earlier observations in Figure 9, $\ell_{1.1}$-MD produces markedly sparser attention weights compared to Adam, with a substantially higher concentration of near-zero parameters. This sparsity pattern reveals that $\ell_{1.1}$-MD implicitly performs feature selection during training, concentrating representational capacity on the most discriminative attention components while driving less relevant weights toward zero. Such induced sparsity not only enhances model interpretability by identifying which attention mechanisms are critical for the task, but also enables more aggressive pruning and compression strategies without sacrificing performance—a practical advantage for deployment in resource-constrained environments.
@@
Similarly, we also compare these algorithms in training the same ViT architecture on CIFAR-100, a more complex dataset.
The results in Figure LABEL:fig:response-sec4-test-acc show that our MD algorithms generally achieve test scores of around $55\%$, which are at a similar level to those of other popular training algorithms, such as SGD and Adam, on CIFAR-100, further demonstrating that our algorithms can compete with the state-of-the-art. Again, we plot histograms of the absolute values of the key, query, and value weights in the attention layers of the resulting models that were trained using Adam and $\ell_{1.1}$MD. Figure 12 shows these histograms, and as we can see, the resulting model trained by $\ell_{1.1}$MD has significantly more attention weight entries that are close to zero when compared to the model trained by Adam, making the attention weights sparser.
Figure 12: Histogram of the absolute values of the entries in the $W_{K},W_{Q}$, and $W_{V}$ attention weight matrices of the resulting models trained by $\ell_{1.1}$MD and Adam on CIFAR-100. We can see that the weights from the model that was trained by $\ell_{1.1}$MD is more sparse. The red dot represent the maximum absolute values of the respective weight matrices.
5 Conclusion and Future Work